# Adaptive Incentive Design with Learning Agents

### v1

@@ 4.1 Atomic Aggregative Games
Here, we study quadratic network aggregative games [BK07, BGR16, SCM22a, AJ13]. Consider a game $G$ comprised of a finite set of players $\mathcal{I}$. The strategy set of every player is the entire real line $\mathbb{R}$. Given the joint strategy profile $x=(x_{i})_{i\in\mathcal{I}}$, the cost of each player $i\in\mathcal{I}$ is
@@
Such separable social costs have been considered in the literature on control of agents in aggregative games [Gra17] as well as systemic risk analysis in financial networks [AOTS15]. One example of $\Phi(x)$ that satisfies this property include $\Phi(x)=\frac{1}{2}\sum_{i\in\mathcal{I}}(x_{i}-\xi_{i})^{2}$, which computes the difference between the players’ strategies and some desired strategy $\xi\in\mathbb{R}^{|\mathcal{I}|}$. Another example is $\Phi(x)=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}x_{i}$ representing the aggregate level of all agents’ activities in the network [AOTS15].
@@ 4.2 Non-atomic Traffic Routing on General Networks
For any o-d pair $i\in\tilde{\mathcal{I}}$ and route flow $\tilde{x}\in\mathbb{R}^{|\tilde{R}|}$, the cost experienced by travelers using route $j\in\tilde{R}_{i}$ is ~$\tilde{\ell}_{i}^{j}(\tilde{x})=\sum_{a\in\tilde{\mathcal{E}}}l_{a}(\tilde{w}_{a})\mathbbm{1}(a\in j)$, where $l_{a}(\cdot)$ is the edge latency function that depends on the edge flows. For every edge $a\in\tilde{\mathcal{E}}$, we assume that the edge latency function $l_{a}(\cdot)$ is convex and strictly increasing. This property of edge latency function captures the congestion effect on the transportation network [BMW56, Rou10].
@@
Consequently, the total cost experienced by travelers on o-d pair $i\in\tilde{\mathcal{I}}$ who choose route $j\in\tilde{R}_{i}$ is $\tilde{c}_{i}^{j}(\tilde{x},\tilde{P})=\tilde{\ell}_{i}^{j}(\tilde{x})+\tilde{P}_{i}^{j}$. Let $\tilde{x}^{\ast}(\tilde{P})$ denote a Nash equilibrium corresponding to route tolls $\tilde{P}$. Owing to (35), with slight abuse of notation, we shall frequently use $\tilde{x}^{\ast}(\tilde{P})$ and $\tilde{x}^{\ast}(\tilde{p})$ interchangeably. Typically, the equilibrium route flows can be non-unique but the corresponding edge flows $\tilde{w}^{\ast}(\tilde{p})$ are unique. Furthermore, the function $\tilde{p}\mapsto\tilde{w}^{\ast}(\tilde{p})$ is a continuous function [YH05].
@@
Note that the optimal route flow can be non-unique but the optimal edge flow, denoted by $w^{\dagger}$, is unique [YH05].
@@
Note that for any arbitrary edge toll $\tilde{p}\in\mathbb{R}^{|\tilde{\mathcal{E}}|}$, $\tilde{w}^{\ast}(\tilde{p})$ is the unique solution to the following strictly convex optimization problem [San10].
@@
Condition (44) ensures that the function $\tilde{\phi}$ is differentiable in a neighborhood of $\tilde{p}^{\dagger}$ (cf. [YH05, Chapter 4]). From Lemma B.1, we know that if $z^{\top}\Delta D\phi(\tilde{p}^{\dagger})z\leqslant 0$ for all $z$ then there exists $r>0$ such that (46) holds for all $\tilde{p}\in\mathcal{B}_{r}(\tilde{p}^{\dagger})$. Indeed, by design of $\Delta$ it holds that $\Delta D\phi(\tilde{p}^{\dagger})=D\tilde{w}(\tilde{p}^{\dagger})$. Finally, using Lemma B.1 and Lemma B.4, we note that $z^{\top}D\tilde{w}(\tilde{p}^{\dagger})z\leqslant 0$. This concludes the proof. ∎
@@ 5 Concluding Remarks
An important direction for future research is to verify the convergence conditions in a broader class of games. Notably, leveraging the externality-based dynamic incentive design update proposed in an earlier version of this paper (cf. [MKWS21]), [LMZ23] studied the convergence of such updates in power systems.
@@
In this game, the strategy set is $\tilde{X}=\{\tilde{x}\in\mathbb{R}^{2}:\tilde{x}_{1}+\tilde{x}_{2}=1\}$. The equilibrium congestion levels on the two edges is obtained by computing the minimizer of the following function [YH05]:
@@
However, this function is not convex, for example, on the set $|\tilde{p}_{1}-\tilde{p}_{2}|<1.5$. Therefore, gradient based updates initialized in the region $\{(\tilde{p}_{1},\tilde{p}_{2})\in\mathbb{R}^{2}:1.1\leqslant|\tilde{p}_{1}-\tilde{p}_{2}|\leqslant 1.5\}$ will not converge to the global optimizer as the gradient is zero in this region. Thus, the approach from [LNL+23, LLY+21] does not apply here.

### v2

@@ 4.1 Atomic Aggregative Games
Here, we study quadratic network aggregative games [Bramoullé and Kranton, 2007, Bramoullé et al., 2016, Shakarami et al., 2022a, Acemoglu and Jensen, 2013]. Consider a game $G$ comprised of a finite set of players $\mathcal{I}$. The strategy set of every player is the entire real line $\mathbb{R}$. Given the joint strategy profile $x=(x_{i})_{i\in\mathcal{I}}$, the cost of each player $i\in\mathcal{I}$ is
@@
Such separable social costs have been considered in the literature on control of agents in aggregative games [Grammatico, 2017] as well as systemic risk analysis in financial networks [Acemoglu et al., 2015]. One example of $\Phi(x)$ that satisfies this property include $\Phi(x)=\frac{1}{2}\sum_{i\in\mathcal{I}}(x_{i}-\xi_{i})^{2}$, which computes the difference between the players’ strategies and some desired strategy $\xi\in\mathbb{R}^{|\mathcal{I}|}$. Another example is $\Phi(x)=\frac{1}{|\mathcal{I}|}\sum_{i\in\mathcal{I}}x_{i}$ representing the aggregate level of all agents’ activities in the network [Acemoglu et al., 2015].
@@ 4.2 Non-atomic Traffic Routing on General Networks
For any o-d pair $i\in\tilde{\mathcal{I}}$ and route flow $\tilde{x}\in\mathbb{R}^{|\tilde{R}|}$, the cost experienced by travelers using route $j\in\tilde{R}_{i}$ is ~$\tilde{\ell}_{i}^{j}(\tilde{x})=\sum_{a\in\tilde{\mathcal{E}}}l_{a}(\tilde{w}_{a})\mathbbm{1}(a\in j)$, where $l_{a}(\cdot)$ is the edge latency function that depends on the edge flows. For every edge $a\in\tilde{\mathcal{E}}$, we assume that the edge latency function $l_{a}(\cdot)$ is convex and strictly increasing. This property of edge latency function captures the congestion effect on the transportation network [Beckmann et al., 1956, Roughgarden, 2010].
@@
Consequently, the total cost experienced by travelers on o-d pair $i\in\tilde{\mathcal{I}}$ who choose route $j\in\tilde{R}_{i}$ is $\tilde{c}_{i}^{j}(\tilde{x},\tilde{P})=\tilde{\ell}_{i}^{j}(\tilde{x})+\tilde{P}_{i}^{j}$. Let $\tilde{x}^{\ast}(\tilde{P})$ denote a Nash equilibrium corresponding to route tolls $\tilde{P}$. Owing to (35), with slight abuse of notation, we shall frequently use $\tilde{x}^{\ast}(\tilde{P})$ and $\tilde{x}^{\ast}(\tilde{p})$ interchangeably. Typically, the equilibrium route flows can be non-unique but the corresponding edge flows $\tilde{w}^{\ast}(\tilde{p})$ are unique. Furthermore, the function $\tilde{p}\mapsto\tilde{w}^{\ast}(\tilde{p})$ is a continuous function [Yang and Huang, 2005].
@@
Note that the optimal route flow can be non-unique but the optimal edge flow, denoted by $w^{\dagger}$, is unique [Yang and Huang, 2005].
@@
Note that for any arbitrary edge toll $\tilde{p}\in\mathbb{R}^{|\tilde{\mathcal{E}}|}$, $\tilde{w}^{\ast}(\tilde{p})$ is the unique solution to the following strictly convex optimization problem [Sandholm, 2010].
@@
Condition (44) ensures that the function $\tilde{\phi}$ is differentiable in a neighborhood of $\tilde{p}^{\dagger}$ (cf. [Yang and Huang, 2005, Chapter 4]). From Lemma B.1, we know that if $z^{\top}\Delta D\phi(\tilde{p}^{\dagger})z\leqslant 0$ for all $z$ then there exists $r>0$ such that (46) holds for all $\tilde{p}\in\mathcal{B}_{r}(\tilde{p}^{\dagger})$. Indeed, by design of $\Delta$ it holds that $\Delta D\phi(\tilde{p}^{\dagger})=D\tilde{w}(\tilde{p}^{\dagger})$. Finally, using Lemma B.1 and Lemma B.4, we note that $z^{\top}D\tilde{w}(\tilde{p}^{\dagger})z\leqslant 0$. This concludes the proof. ∎
@@ 5 Concluding Remarks
An important direction for future research is to verify the convergence conditions in a broader class of games. Notably, leveraging the externality-based dynamic incentive design update proposed in an earlier version of this paper (cf. [Maheshwari et al., 2021]), [Li et al., 2023a] studied the convergence of such updates in power systems.
Acknowledgements
This work is supported by NSF Collaborative Research: Transferable, Hierarchical, Expressive, Optimal, Robust, Interpretable NETworks (THEORINET) under award No. DMS-2031899.
@@ Acknowledgements
In this game, the strategy set is $\tilde{X}=\{\tilde{x}\in\mathbb{R}^{2}:\tilde{x}_{1}+\tilde{x}_{2}=1\}$. The equilibrium congestion levels on the two edges is obtained by computing the minimizer of the following function [Yang and Huang, 2005]:
@@
However, this function is not convex, for example, on the set $|\tilde{p}_{1}-\tilde{p}_{2}|<1.5$. Therefore, gradient based updates initialized in the region $\{(\tilde{p}_{1},\tilde{p}_{2})\in\mathbb{R}^{2}:1.1\leqslant|\tilde{p}_{1}-\tilde{p}_{2}|\leqslant 1.5\}$ will not converge to the global optimizer as the gradient is zero in this region. Thus, the approach from [Li et al., 2023b, Liu et al., 2021] does not apply here.